### v1

@@ 3.2 Sharpness for Gaussian Processes
In the setting of Theorem 1, suppose that $(X_{t})_{t\in T}$ is a Gaussian process (with metric defined in (8)). For any $\mu$, we have
$\mathbb{E}[X_{Z}]\geq c\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (36)
where the supremum is over all $Z(\omega)$ satisfying $P_{Z}=\mu$, and
$c:=\max_{\tau\in(0,4/5)}\left\{\frac{4-5\tau}{2\sqrt{2}(4-\tau)\ln\frac{4}{\tau}}\right\}\approx 0.0935918,$ (37)
numerically evaluated at $\tau=0.23457905$.
Theorem 3 suggests that Theorem 1 may yield sharper results in certain statistical applications where the traditional method based on Dudley’s integral is suboptimal. Three proofs of Theorem 3 are given in Section 5: First, the result can be proved using the lifting argument, by applying the majorizing measure theorem to the $\mu$-type set. Secondly, we can apply the lifting argument and use Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes, i.e., the case where the metric on $T$ is invariant under the action of a transitive group (see (Dudley,, 1999, Cor 2.7.5) or (Fernique,, 1975)), which preceded Talagrand’s striking proof of the majorizing measure theorem (Talagrand,, 1987). Even though $T$ in Theorem 3 may not be permutation-invariant, the lifting argument in our proof lower bounds the expected supremum over the typical set, which satisfies permutation invariance. The third proof derives an information-theoretic counterpart of the super-Sudakov inequality (Talagrand,, 1992; Van Handel,, 2014), and finishes off by purely analytic arguments (without chaining or packing constructions), which gives the constant stated in Theorem 3.
Classically, equivalent formulations of the majorizing measure theorem, (14)-(15), were established by Fernique convexity argument; see (Talagrand,, 2014, Chapter 3) for a proof using combinatorial arguments such as tree packing. Similarly, the rate-distortion integral can also be expressed in alternative forms by applying the minimax theorem to a certain convex-concave functional; see Remark 1 for discussions.
We can derive from Theorem 3 the classical the majorizing measure theorem, with an explicit estimate of constant. Recall $I_{\mu}$ defined in (11).
@@
For any Gaussian process (not necessarily stationary),
$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (38)
Here $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{\frac{2a}{c}+\sqrt{2\pi h(a^{-2})}}\approx 0.0218988$ (evaluated at $a=1.4392$), where $h$ is the binary entropy function in nats and $c$ was defined in Theorem 3.
The proof of Corollary 1 is given in Section 6, which is essentially based on lower bounding the rate-distortion function by the probability of a set using the data processing inequality. We have not found a reference on the majoring measure theorem that estimated the constant to compare with Corollary 1.
@@
The proof immediately follows from Theorem 1-3, since the rate-distortion functions under the two metrics can be directly compared. Corollary 2 can be viewed as a “soft” version of (13) (in fact, it is easy to see that it implies (13)). The Euclidean special case of Corollary 2 is of particular interest in practice:
@@
Although Corollary 3 looks like a simple Cauchy-Schwarz inequality, it would not be true if sub-Gaussianity is replaced by a second-moment bound. More precisely, “the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ is dominated by the Euclidean metric” cannot be replaced by $(X_{t})$ is a stochastic process satisfying $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$”. Indeed, consider $T$ as a maximal $\epsilon$-packing of the unit ball in $\mathbb{R}^{N}$, where $N=3$, and suppose that $X_{t}$ is an independent zero mean random variable with $\epsilon^{2}$ variance for each $t\in T$. Then it is true that $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$. However, for fixed $N=3$ and vanishing $\epsilon$, it is possible that $\mathbb{E}[X_{Z}]=\mathbb{E}[\max_{t\in T}X_{t}]\asymp\epsilon\sqrt{|T|}\asymp\epsilon^{1-N/2}$ diverges, for certain heavy-tailed $X_{t}$. On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})=O(1)$ as $\epsilon$ vanishes, confirming that this is a counterexample.
3.3 Tail Bounds for ERM
In this section we consider processes in the form of an inner product in $\ell_{2}$. In this case, the complexity of an index distribution $\mu$ (Definition 3) is written as
$\mathsf{w}(\mu):=\sup_{P_{YZ}\in\Pi(\mu,\gamma)}\mathbb{E}[\left<Y,Z\right$>], (39)
where $\gamma$ is a given distribution on $\ell_{2}$, which is assumed to have zero mean in this section. We may denote $\mathsf{w}(\mu)$ by $\mathsf{w}(\mu;\gamma)$ to make the dependence on $\gamma$ explicit. The following regularity condition will be used in this section:
Definition 4.
$T\subseteq\ell_{2}$ is totally bounded if the covering number $N(T,\epsilon)$ is finite for every $\epsilon>0$.
In empirical process theory, total boundedness is a rather mild condition which is almost always assumed; see for example Ledoux and Talagrand, (2013). It is sometimes also referred to as precompactness in topology. It is also equivalent to
$\lim_{n\to\infty}\sup_{t\in T}\sum_{i>n}t(i)^{2}=0.$ (40)
Indeed, if (40) is true, then for every $\epsilon>0$ there exists $n$ such that $\sup_{t\in T}\sum_{i>n}t(i)^{2}<(\epsilon/4)^{2}$. We can cover the projection of $T$ onto the first $n$ coordinates with finitely many balls of radius $\epsilon/2$, whose centers are the projections of $t_{1},\dots,t_{M}\in T$. Then $B(t_{1},\epsilon),\dots,B(t_{M},\epsilon)$ also cover $T$. Conversely, total boundedness implies that every sequence in $T$ has a convergent subsequence, which implies (40).
Definition 5.
Suppose that $T$ is a subset of $\ell_{2}$. We call $\mathsf{G}\colon[0,\infty)\to[0,\infty)$ a width-moment function of $T$ if for any random vector $Z\in T$ satisfying $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$, we have
$\mathsf{w}(P_{Z})\leq\mathsf{G}(E).$ (41)
Consider the setting of the empirical risk minimization problem in Section 2.2 (either the mean estimation or the regression problem). Our main result, for $\gamma$ satisfying the log-Sobolev inequality, is the following:
@@ 3.3 Tail Bounds for ERM
Suppose that $m\in T$, where $T$ is a closed, totally bounded subset of $\ell_{2}$, and TTT-T has a width-moment function $\mathsf{G}$. Assume that $\gamma$ has zero mean and satisfies the log-Sobolev inequality (17). Let $\psi(\lambda):=\ln\mathbb{E}[\exp(\lambda\|m-\hat{m}\|_{2}^{2})]$. Then for any $\lambda\geq 0$,
$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\epsilon+\frac{2\lambda}{\sqrt{n}}\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\}.$ (42)
The proof of Theorem 4 is given in Section 7. The log-Sobolev inequality assumption in Theorem 4 can in fact be weakened to sub-Gaussianity. In that case, define
$\mathsf{w}_{\rm g}(\mu):=\sup_{P_{YZ}\in\Pi(\gamma_{\rm g},\mu)}\mathbb{E}\left[\sum_{i=1}^{\infty}Y(i)Z(i)\right],$ (43)
where $\gamma_{\rm g}$ denotes the distribution of an infinite sequence of i.i.d. standard normal random variables, and the infinite sum is defined in terms of convergence in second moment. Similarly, define $\mathsf{G}_{\rm g}(E)$ by replacing the left side of (41) with $\mathsf{w}_{\rm g}(P_{Z})$. Note that if $\gamma$ satisfies the log-Sobolev inequality (hence also sub-Gaussian), then Corollary 2 implies that $\mathsf{w}_{\rm g}$ and $\mathsf{G}_{\rm g}$ dominate $\mathsf{w}$ and $\mathsf{G}$ respectively. We have the following result:
Theorem 5.
In the setting of Theorem 4, suppose that $\gamma$ is only assumed to have zero-mean and satisfy $\|Y\|_{\Psi_{2}}\leq 1$, $Y\sim\gamma$. Assume that $\mathsf{G}_{\rm g}(\cdot)$ is continuous. Then for any $\lambda\geq 0$,
$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\frac{\epsilon}{C}+\frac{2K^{\prime}C\lambda}{\sqrt{n}}\mathsf{G}_{\rm g}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+K^{\prime}C\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\},$ (44)
where $C,K^{\prime}>0$ are universal constants in Lemma 2 and (10), respectively.
The proof of Theorem 5 is given in Section 7.2. Similar to the remark following Lemma 2, Theorem 5 assumes a weaker condition on $\gamma$ than (4), but may yield a weaker bound if $\mathsf{G}_{\rm g}$ is larger than $\mathsf{G}$ (for example, in the case where $Y$ has i.i.d. coordinates following the uniform distribution on [1,1][-1,1]).
3.4 ERM for the Sobolev Ellipsoid
We apply the general technique in Section 3.3 to Sobolev ellipsoids, which plays a prominent role in nonparametric statistics (Van de Geer,, 1990; Wainwright,, 2019; Tsybakov,, 2009). The precise problem of estimation over a Sobolev ellipsoid was previously studied in Wei et al., (2020). Generally, an ellipsoid in $\ell_{2}$ is defined as
where $a_{1},a_{2},\dots$ is a decreasing sequence of nonnegative numbers. This is recurring example in Talagrand, (2014), in particular an example showing that Dudley’s integral is loose. A Sobolev ellipsoid is an ellipsoid with $a_{i}=i^{-\beta}$ for some $\beta>0$. By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class (Tsybakov,, 2009). In Section 9.1, we will review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
Consider ERM in the setting of Section 2.2, where $T=\phi(\mathcal{E})$ is the image of the Sobolev ellipsoid (for $\beta>1/2$, where $1/2$ is the critical exponent for ERM) under a 1-Lipschitz map $\phi\colon\ell_{2}\to\ell_{2}$. The purpose of introducing $\phi$ is that if $\phi$ is linear, an elementary proof exists, while the rate-distortion integral approach can handle more general settings. Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$. Under these assumptions, we have the following main result:
Theorem 6.
@@ 3.4 ERM for the Sobolev Ellipsoid
There exists a function $c(\beta)>0$ satisfying $c(\beta)=O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow\frac{1}{2}$, and a universal constant $b>0$ such that
$\mathbb{P}[\|m-\hat{m}\|_{2}>\sqrt{c(\beta)}n^{-\frac{\beta}{2\beta+1}}]\leq\exp(-bc(\beta)n^{\frac{1}{2\beta+1}}).$ (46)
@@
Theorem 6 shows that ERM achieves square error $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ with high probability, which is optimal not only in terms of the dependence on $n$, but also in terms of the behavior of the preconstant $O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow 1/2$. Furthermore, from the proof of Theorem 6 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$.
We present two approaches towards Theorem 6 in Section 9.2. The first one does not use any new tools introduced in this paper, but only works for a special case of Theorem 6 where $\phi$ is the identity map and $Y$ satisfies the stronger log-Sobolev inequality: We use the Cauchy-Schwarz argument to bound the mean of the square error, and then use a concentration inequality to control the deviation from the mean. This approach is a slightly streamlined version of the method in Chatterjee, (2014); Wei et al., (2020). The second proof follows the general recipe in Section 3.3, and works for general sub-Gaussian $Y$ and nonlinear Lipschitz $\phi$, where the rate-distortion integral is used to bound $\mathsf{G}_{\rm g}$.
For comparison, the traditional techniques will only prove an upper bound at the rate $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$, which is optimal in terms of the dependence on $n$, but not in terms of the dependence on $\beta$ in the preconstant as $\beta\downarrow 1/2$ (we analyze the preconstant using the traditional method in Remark 3). Our result shows that the optimal $O(\frac{1}{\sqrt{2\beta-1}})$ preconstant can be regained by using the rate-distortion integral instead.
3.5 ERM for Sparse Regression
This section provides new upper bounds for empirical risk minimization for the weak $\ell_{q}$ ball in $\mathbb{R}^{d}$ defined in (4), extending prior results for the $\ell_{q}$ ball (3). Consider a special case of the general regression model in Section 2.2, where $T$ is chosen as a linear transform of $wB_{q}(r)$. More precisely, suppose that the columns of $\frac{1}{\sqrt{n}}\mathbb{X}$ have unit $\ell_{2}$ norm, $\beta^{*}\in wB_{q}(r)$, $q\in(0,1)$, and $Y=\mathbb{X}\beta^{*}+w$, where $w$ is a zero mean random vector satisfying $\|w\|_{\Psi_{2}}\leq 1$. Let $\hat{\beta}:=\operatorname*{arg\,min}_{\beta\in wB_{q}(r)}\|Y-\mathbb{X}\beta\|_{2}$. Throughout, we set
$R:=r^{q},$ (47)
and we caution the reader that some prior work, including Raskutti et al., (2011), uses the notation $B_{q}(R)=B_{q}(r^{q})$ to refer to what we denote by $B_{q}(r)$. Applying the technique in Section 3.3, we obtain the following result:
Theorem 7.
There exists $C_{q}>0$ (depending only on $q\in(0,1)$) such that
$\mathbb{P}\left[\|m-\hat{m}\|_{2}^{2}>C_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}R\right]\leq\exp(-(\ln d)^{\frac{2-q}{2}}Rn^{\frac{q}{2}}).$ (48)
The proof of Section 10 is given in Theorem 7, which is based on the rate-distortion integral and the general technique for tail bounds in Section 3.3.
Previously, Raskutti et al., (2011) established the sharp convergence of ERM when $\beta\in B_{q}(r)$ where $q\in(0,1]$. The analysis for the case of $\beta\in B_{1}(r)$ follows by a simple Hölder inequality argument (see Raskutti et al., (2011)), whereas the $q<1$ case was proved by the slicing technique, based on the Dudley integral where the covering numbers were bounded using deep results from the convex geometry (Carl and Pajor,, 1988; Guedon and Litvak,, 2000). In contrast, we prove Theorem 7 using the rate-distortion integral, where the rate-distortion function (a priori smaller than the log covering number, hence easier to upper bound) is upper bounded using purely information-theoretic tools, without invoking the convex geometry literature. Furthermore, Theorem 7 shows that the minimax rate for $wB_{q}(r)$ is the same as $B_{q}(r)$ when $q\in(0,1)$, a nontrivial extension of the result of Raskutti et al., (2011) which appears to be beyond the previous techniques. The $q=1$ case is somewhat special: by considering $\mathbb{X}$ with identical rows, we can see that the rate for $wB_{1}(r)$ can be worse than that of $B_{1}(r)$ by a logarithmic factor. On the other hand, in the special case where $\frac{1}{\sqrt{n}}\mathbb{X}$ is the identity, the rate $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ is achievable for $wB_{q}(r)$ for all $q\in(0,2)$; see Rigollet and Hütter, (2023); Donoho and Johnstone, (1994); Zhang, (2010).
4 Proof of the Rate-Distortion Integral
In this section we prove Theorem 1, Theorem 2, and Proposition 1.
4.1 Proof in the Finite Case
Recall the following basic property of the sub-Gaussian distribution:

### v3

@@ 3.2 Sharpness for Gaussian processes
Suppose that $(X_{t})_{t\in T}$ is a centered Gaussian process, $T$ is a finite set, and $\mu$ is an arbitrary distribution on $T$. Then have
$K^{\prime}\mathsf{w}(\mu)\geq\int_{0}^{\infty}\sqrt{i_{\mu}(\sigma)}d\sigma\geq\int_{0}^{\sigma_{\rm m}(\mu)}\sqrt{R_{\mu}(\sigma^{2})}d\sigma,$ (34)
where $K^{\prime}$ is any universal constant such that (7) holds.
We present a proof of Theorem 3 in Section 5, by using Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes (even though Theorem 3 does not assume stationary process). In the appendix we also extend the result to the case of countable $T$. Theorem 3 suggests that Theorem 1 may yield sharper results in certain statistical applications where the traditional method based on Dudley’s integral is suboptimal. From Theorem 3 we can derive the classical majorizing measure theorem. Recall $I_{\mu}$ defined in (8).
@@
For any Gaussian process (not necessarily stationary), we have
$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t),$ (35)
where $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{2aK^{\prime}+\sqrt{2\pi h(a^{-2})}}$, $h(\cdot)$ is the binary entropy function, and $K^{\prime}$ is any universal constant such that (34) holds.
The proof of Corollary 1 is given in Section 6, which is essentially based on lower bounding the rate-distortion function by the probability of a set using the data processing inequality.
@@
The proof immediately follows from Theorem 1-3, since the rate-distortion functions under the two metrics can be directly compared. Corollary 2 can be viewed as a “soft” version of (10) (in fact, it is easy to see that it implies (10)). The Euclidean special case of Corollary 2 is of particular interest and will be used later in our statistical applications:
@@
Although Corollary 3 looks like a simple Cauchy-Schwarz inequality, it would not be true if sub-Gaussianity is replaced by a second-moment bound. More precisely, “the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ is dominated by the Euclidean metric” cannot be replaced by $(X_{t})$ is a stochastic process satisfying $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$”. Indeed, consider $T$ as a maximal $\epsilon$-packing of the unit ball in $\mathbb{R}^{N}$, where $N=3$, and suppose that $X_{t}$ is an independent zero mean random variable with $\epsilon^{2}$ variance for each $t\in T$. Then it is true that $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$. However, for fixed $N=3$ and vanishing $\epsilon$, it is possible that $\mathbb{E}[X_{Z}]=\mathbb{E}[\max_{t\in T}X_{t}]\asymp\epsilon\sqrt{|T|}\asymp\epsilon^{1-N/2}$ diverges, for certain heavy-tailed $X_{t}$. On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})$ remains bounded from above as $\epsilon$ vanishes, confirming that this is a counterexample.
3.3 Application to ERM over the Sobolev ellipsoid
We use the rate-distortion integral to upper bound the error in ERM over a Sobolev ellipsoid under a Lipschitz transformation, A general ellipsoid in $\ell_{2}$ is defined as
where $a_{1},a_{2},\dots$ is a decreasing sequence of nonnegative numbers. A Sobolev ellipsoid is an ellipsoid with $a_{i}=i^{-\beta}$ for some $\beta>0$. By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class [40]. Sobolev ellipsoids therefore play a prominent role in nonparametric statistics [41, 44, 40]. The precise problem of ERM over a Sobolev ellipsoid was previously studied in [45].
The ellipsoid is recurring example in [39], in particular an example showing that Dudley’s integral is loose. In Section 7, we review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
Consider ERM in the setting of Section 2.2, where $T=\phi(\mathcal{E})$ is the image of the Sobolev ellipsoid (for $\beta>1/2$, where $1/2$ is the critical exponent for ERM) under a 1-Lipschitz map $\phi\colon\ell_{2}\to\ell_{2}$. In practice, $\phi$ may represent a nonlinear neural network or feature map. If $\phi$ is the identity map, the Cauchy-Schwarz inequality suffices for the error upper bound (Remark 2), while the rate-distortion integral machinery is essential for general $\phi$, analogous to the role of the majorizing measure theorem for sub-Gaussian comparison (10). Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{G}\leq 1$. Under these assumptions, we have the following main result:
@@ 3.3 Application to ERM over the Sobolev ellipsoid
There exists a function $c(\beta)>0$ satisfying $\lim_{\beta\downarrow\frac{1}{2}}\sqrt{2\beta-1}c(\beta)<\infty$, such that $\mathbb{E}[\|m-\hat{m}\|_{2}^{2}]\leq c(\beta)n^{-\frac{2\beta}{2\beta+1}}$.
We present a proof of Theorem 4 in Section 7.3.
Remark 1.
@@
Theorem 4 shows that ERM achieves mean squared error of order $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$ with high probability. Furthermore, from the proof of Theorem 4 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$. For comparison, the traditional technique of [41] based on Dudley’s integral will only prove an upper bound at the rate $\frac{1}{2\beta-1}n^{-\frac{2\beta}{2\beta+1}}$ (see Remark 4), which has the optimal exponent of $n$, but a worse preconstant.
Remark 2.
For convex regression classes with Gaussian noise, [7] showed that the mean-squared error (MSE) of ERM can be characterized by a fixed-point equation involving a localized Gaussian width. For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality [45], thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound. However, the Cauchy-Schwarz step of [45] exploits linear structure, and cannot be applied to the setting of Theorem 4 when the map $\phi$ is nonlinear Lipschitz. This limitation is expected to be artifacts of the proof technique of [7, 45]: Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (10)). Our work thus introduces a new proof technique that overcomes this limitation.
4 Proof of the rate-distortion integral: finite case
In this section we prove Theorem 1 and Theorem 2 in the case of finite index set; extension beyond the finite case follows by a limiting argument, and is deferred to the appendix. Recall the following basic property of the sub-Gaussian distribution, whose proof can be found in [43, Exercise 2.40]: